OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 2, -4, -10, ? -130 -124 -118 -112

OpenStudy (anonymous):

\[a_n = a+ (n-1)d\] a_n is the n'th term d-->common difference a-->first term

jimthompson5910 (jim_thompson5910):

a[n] = a[1] + (n-1)*d a[n] = 2 + (n-1)*(-6) a[n] = 2 -6n + 6 a[n] = -6n + 8 a[20] = -6(20) + 8 a[20] = -112 So the 20th term is -112

OpenStudy (anonymous):

So in your problem, n=20,you know, a, d . Find a_n